OpenStudy (anonymous):

Prove or give a counter example:

OpenStudy (anonymous):

If \[\sum_{n=1}^{\infty} a_n\] and \[\sum_{n=1}^{\infty} b_n\] are both convergent series with positive terms, then \[\sum_{n=1}^{\infty} b_n * a_n\] is also convergent.

OpenStudy (anonymous):

I think the property: If the series of a_n is convergent then lim as n approaches infinity of a_n =0 will help.

OpenStudy (anonymous):

I'm stuck on where to go next and it would be terrific if someone could help :)

OpenStudy (anonymous):

My teacher gave me a hint and said when lim n→∞ bn = 0 it means that for large enough n, say n > N for some large N, bn ≤ 1? I don't understand this..

OpenStudy (rational):

Ahh nice that will be a good start

OpenStudy (anonymous):

I'm not sure that I understand it though. how does b_n >= 1?

OpenStudy (rational):

\[\sum\limits_{n=1}^{\infty}b_n\] if this series converges, then the terms \(b_n\) approach \(0\) as \(n\) gets large, yes ?

OpenStudy (anonymous):

yes

OpenStudy (rational):

that means there exists some index position \(N\), after which all the terms \(b_n\) are less than \(1\), yes ?

OpenStudy (anonymous):

Yes

OpenStudy (rational):

thats what the hint says

OpenStudy (anonymous):

Okay... so can we split the series in that sort then?

OpenStudy (rational):

Yes

OpenStudy (rational):

\(b_n\le 1 \implies a_nb_n\le a_n\) we may use comparison test

OpenStudy (anonymous):

"Note that since the series of b_n from 1 to infinity is convergent, the lim of b_n as n approaches infinity =0. This means that if n is large enough there exists some index point N after all the terms in the sequence b_n are less than 1.

OpenStudy (anonymous):

Would that be a valid statement to write on my homework??

OpenStudy (anonymous):

Why do we say n> N?

OpenStudy (rational):

lets consider a quick example

OpenStudy (anonymous):

Awesome! thank you :)

OpenStudy (rational):

let \(b_n=\frac{10}{n^2}\) can you plot first few terms of \(\sum\limits_{n=1}^{\infty}b_n\) on number line?

OpenStudy (anonymous):

sure thang

OpenStudy (anonymous):



OpenStudy (anonymous):

right to left, a_1, a_2, a_3, a_4, and a_5

OpenStudy (rational):